# Fundamental limits of weak learnability in high-dimensional multi-index models

### v1

@@ 5 Hierarchical Iterative Denoising
almost surely over ${\bm{\xi}}\sim\mathcal{N}(\mathbf{0},{\bm{I}}_{p})$ and $Y$ for any ${{\bm{M}}}_{U}\in\mathcal{S}^{+}_{p}$ such that $\operatorname{span}({{\bm{M}}}_{U})=U$. We define the “trivially-denoising-coupled” subspace $T^{\star}_{U}$ for $U$ as the orthogonal complement of $H_{T}^{\star}(U)$.
@@
almost surely over $bxi$ and $Y$ for any ${{\bm{M}}}_{U}\in\mathcal{S}^{+}_{p}$ such that $\operatorname{span}({{\bm{M}}}_{U})=U$.
@@
Let $U\in\mathbb{R}^{p}$ be a fixed subspace. Suppose the AMP algorithm 1 is initialised such that ${{\bm{M}}}_{d}^{0}={{\bm{M}}}_{U}+\epsilon{\bm{A}}$, where $\norm{{\bm{A}}}=1$ and ${{\bm{M}}}_{U}$ is a fixed point of $F({{\bm{M}}})$ in Equation 8 with $\operatorname{span}({{\bm{M}}}_{U})=U$. If $\alpha\geq\alpha_{\text{hid}}({{\bm{M}}}_{U})$, $\exists\delta>0$ such that for any sufficiently small $\epsilon$ ,${{\bm{M}}}^{(t)}\succ\delta{{\bm{M}}}^{*}_{U}$ for some $t=\mathcal{O}(\log\nicefrac{{1}}{{\epsilon}})$, where ${{\bm{M}}}^{*}_{U}$ is defined as in Definition 5. For $\alpha<\alpha_{\text{hid}}({{\bm{M}}}_{U})$ however, there exists an $\epsilon^{\prime}$ such that for $\epsilon<\epsilon^{\prime}$, $\norm{{{\bm{M}}}^{t}_{\perp}}=0$ as $t\rightarrow\infty$, where ${{\bm{M}}}^{t}_{\perp}$ denotes the projection of ${{\bm{M}}}^{t}$ orthogonal to $U$. Furthermore, suppose that ${\bm{A}}$ is full-rank, then there exists an $\alpha>0$ and a $\delta>0$ such that is ${{\bm{M}}}_{d}^{t}\succ\delta{{\bm{M}}}_{E^{\star}_{U}}$, where ${{\bm{M}}}_{E^{\star}_{U}}\in\mathcal{S}^{+}_{p}$ spans $E^{\star}_{U}$
@@
$g({\bm{z}})=z_{1}^{4}+{\rm sign}(z_{1}z_{2}z_{3})+{\rm sign}(z_{3}z_{4}z_{5})+{\rm sign}(z_{5}z_{6}z_{7})+\ldots$ (25)
@@
Finally, it is important to stress that hierarchical learnable functions are not equivalent to staircase functions (Abbe et al., 2022; 2023). For instance, (25) is learned with $\Theta(d)$, while it requires $O(d^{3})$ samples in the CSQ staircase characterisation.
@@ 6 Conclusion and limitations
This work investigated the computational boundaries of learning high-dimensional Gaussian multi-index models with first-order iterative algorithms. As we hinted in the text (and provided some partial answers), a natural question is how close neural networks are to achieving this optimal performance. Two limitations of our work are: (a) our results hold asymptotically in $d\to\infty$; (b) We assume Gaussian covariates. For the former, numerical experiments support that the predictions are already accurate at fairly moderate sizes. For the latter, a number of universality results justify the scope of high-dimensional Gaussian covariates and extend AMP to more structured designs (Loureiro et al., 2021; Wang et al., 2022; Dudeja et al., 2023; Dandi et al., 2024a).
@@ Acknowledgements
The authors would like to thank Joan Bruna, Theodor Misiakiewicz, Luca Pesce, and Nati Srebro for stimulating discussions. BL acknowledges funding from the Choose France - CNRS AI Rising Talents program. This work was supported by the Swiss National Science Foundation under grants SNSF SMArtNet (grant number 212049) and SNSF OperaGOST (grant number 200390).
@@ Appendix A Optimality of AMP
Let $(x_{i},y_{i})_{i\in[n]}$ denote $n$ i.i.d. samples from the multi-index model defined in 1. Denote by $\hat{{{\bm{{W}}}}}_{\rm bo}=\mathbb{E}[W|X,y]\in\mathbb{R}^{p\times d}$ the mean of the posterior marginals Equation 5. Then, in the high-dimensional asymptotic limit where $n,d\to\infty$ with fixed ratio $\alpha=\nicefrac{{n}}{{d}}$, the asymptotic correlation between the posterior mean and ${{\bm{{W}}}}^{\star}$:
@@
Finally, note that by construction $\hat{{{\bm{{W}}}}}_{\rm bo}$ is the optimal estimator of ${{\bm{{W}}}}^{\star}$ given the data $(X,y)$ (in the MMSE sense). Therefore, the rank of $M^{\star}$ defines the dimension of the statistically optimal subspace reconstruction at sample complexity $\alpha\coloneqq\nicefrac{{n}}{{d}}$. The fact that AMP follows state evolution is proven for such problems in Gerbelot and Berthier [2023]
@@ B.1 Linear Approximation
Throughout, we assume that ${\bm{g}}_{\rm out}:\mathbb{R}^{k^{2}+k+1}\rightarrow\mathbb{R}^{k}$ is in $\mathcal{C}^{2}$.
@@
We proceed through an-entry-wise expansion of each term inside the expection in $F({{\bm{M}}})$ around ${{\bm{M}}}=0$. Since ${{\bm{M}}}_{F}\leq p$ , the first two derivatives of ${\bm{g}}_{\rm out}$ are uniformly bounded in ${\bm{g}}_{\rm out}$ for any fixed $Y,{\bm{\xi}}$. Therefore, applying the multivariate Taylor expansion to ${\bm{g}}_{\rm out}(Y^{t},\Delta_{1}{\bm{\xi}},I+\Delta_{2})$, with $\Delta_{1}=\sqrt{M}$ and $\Delta_{2}=,{\bm{I}}_{p}-{{\bm{M}}}$ yields:
@@
$\nu_{\mathcal{F}}=\inf{\mathbf{v}\in\mathbb{E},\norm{v}=1}\mathbf{v}^{\top}\mathcal{F}(\mathbf{v}\mathbf{v}^{\top})v.$ (40)
@@
Therefore, for $\alpha\geq\frac{1}{\nu_{\mathcal{F}}}$ and small enough $\epsilon$, ${{\bm{M}}}^{t}$ expands linearly along each $\mathbf{v}_{i}$
B.4 Proof of Theorem 5.1

### v2

@@ 5 Hierarchical Iterative Denoising
almost surely over ${\bm{\xi}}\sim\mathcal{N}(\mathbf{0},{\bm{I}}_{p})$ and $Y$ for any ${{\bm{M}}}_{U}\in\mathcal{S}^{+}_{p}$ such that $\operatorname{span}({{\bm{M}}}_{U})=U$. We define the “trivially-denoising-coupled” subspace $T^{\star}_{U}$ for $U$ as the orthogonal complement of $H_{E}^{\star}(U)$.
@@
almost surely over ${\bm{\xi}}$ and $Y$ for any ${{\bm{M}}}_{U}\in\mathcal{S}^{+}_{p}$ such that $\operatorname{span}({{\bm{M}}}_{U})=U$.
@@
Let $U\in\mathbb{R}^{p}$ be a fixed subspace. Suppose the AMP algorithm 1 is initialised such that ${{\bm{M}}}_{d}^{0}={{\bm{M}}}_{U}+\epsilon{\bm{A}}$, where $\norm{{\bm{A}}}_{F}=1$ and ${{\bm{M}}}_{U}$ is a fixed point of $F({{\bm{M}}})$ in Equation 7 with $\operatorname{span}({{\bm{M}}}_{U})=U$. Then, under Assumption 1, the following holds with high probability as $d\rightarrow\infty$: If $\alpha\geq\alpha_{\text{hid}}({{\bm{M}}}_{U})$, $\exists\delta>0$ such that for any sufficiently small $\epsilon$ ,${{\bm{M}}}^{(t)}\succ\delta{{\bm{M}}}^{*}_{U}$ for some $t=\mathcal{O}(\log\nicefrac{{1}}{{\epsilon}})$, where ${{\bm{M}}}^{*}_{U}$ is defined as in Definition 5. For $\alpha<\alpha_{\text{hid}}({{\bm{M}}}_{U})$ however, there exists an $\epsilon^{\prime}$ such that for $\epsilon<\epsilon^{\prime}$, $\norm{{{\bm{M}}}^{t}_{\perp}}=0$ as $t\rightarrow\infty$, where ${{\bm{M}}}^{t}_{\perp}$ denotes the projection of ${{\bm{M}}}^{t}$ orthogonal to $U$. Furthermore, suppose that ${\bm{A}}$ is full-rank, then there exists an $\alpha>0$ and a $\delta>0$ such that is ${{\bm{M}}}_{d}^{t}\succ\delta{{\bm{M}}}_{E^{\star}_{U}}$, where ${{\bm{M}}}_{E^{\star}_{U}}\in\mathcal{S}^{+}_{p}$ spans $E^{\star}_{U}$
@@
$g({\bm{z}})=z_{1}^{2}+{\rm sign}(z_{1}z_{2}z_{3})+{\rm sign}(z_{3}z_{4}z_{5})+{\rm sign}(z_{5}z_{6}z_{7})+\ldots$ (26)
@@
Finally, it is important to stress that hierarchical learnable functions are not equivalent to staircase functions (Abbe et al., 2022; 2023). For instance, consider the target:
$g({\bm{z}})=\rm He_{4}(z_{1})+{\rm sign}(z_{1}z_{2}z_{3}).$ (27)
Similar to Example 26, the above target is learned sequentially with $\mathcal{O}(d)$ samples by first recovering the direction corresponding to $z_{1}$. However, under the CSQ staircase classification, $g({\bm{z}})$ corresponds to leap $3$. The CSQ lower-bound therefore implies a sample complexity-requirement of $\Omega(d^{1.5})$ with online-SGD learning it in $\mathcal{O}(d^{2})$ samples (Abbe et al., 2023).
@@ 6 Conclusion and limitations
This work investigated the computational boundaries of learning high-dimensional Gaussian multi-index models with first-order iterative algorithms. As we hinted in the text (and provided some partial answers), a natural question is how close neural networks are to achieving this optimal performance. Two limitations of our work are: (a) our results hold asymptotically in $d\to\infty$; (b) We assume Gaussian covariates. For the former, numerical experiments support that the predictions are already accurate at fairly moderate sizes. For the latter, a number of universality results justify the scope of high-dimensional Gaussian covariates and extend AMP to more structured designs (Loureiro et al., 2021; Wang et al., 2022; Dudeja et al., 2023; Dandi et al., 2023).
@@ Acknowledgements
The authors would like to thank Joan Bruna, Theodor Misiakiewicz, Luca Pesce, and Nati Srebro for stimulating discussions. This work was supported by the Swiss National Science Foundation under grants SNSF SMArtNet (grant number 212049) and SNSF OperaGOST (grant number 200390) and the Choose France - CNRS AI Rising Talents program.
@@ Appendix A Optimality of AMP
Let $(x_{i},y_{i})_{i\in[n]}$ denote $n$ i.i.d. samples from the multi-index model defined in 1. Denote by $\hat{{{\bm{{W}}}}}_{\rm bo}=\mathbb{E}[W|X,y]\in\mathbb{R}^{p\times d}$ the mean of the posterior marginals Equation 5. Then, under Assumption 1 in the high-dimensional asymptotic limit where $n,d\to\infty$ with fixed ratio $\alpha=\nicefrac{{n}}{{d}}$, the asymptotic correlation between the posterior mean and ${{\bm{{W}}}}^{\star}$:
@@
Finally, note that by construction $\hat{{{\bm{{W}}}}}_{\rm bo}$ is the optimal estimator of ${{\bm{{W}}}}^{\star}$ given the data $(X,y)$ (in the MMSE sense). Therefore, the rank of $M^{\star}$ defines the dimension of the statistically optimal subspace reconstruction at sample complexity $\alpha\coloneqq\nicefrac{{n}}{{d}}$. The fact that AMP follows state evolution is proven for such problems in [Javanmard and Montanari, 2013, Gerbelot and Berthier, 2023].
@@ B.1 Linear Approximation
We proceed through an-entry-wise expansion of each term inside the expection in $F({{\bm{M}}})$ around ${{\bm{M}}}=0$. Recall that ${\bm{g}}_{\rm out}\in\mathcal{C}^{2}$ by Assumption 1. Since ${{\bm{M}}}_{F}\leq p$ , the first two derivatives of ${\bm{g}}_{\rm out}$ are uniformly bounded in ${\bm{g}}_{\rm out}$ for any fixed $Y,{\bm{\xi}}$. Therefore, applying the multivariate Taylor expansion to ${\bm{g}}_{\rm out}(Y^{t},\Delta_{1}{\bm{\xi}},I+\Delta_{2})$, with $\Delta_{1}=\sqrt{M}$ and $\Delta_{2}=,{\bm{I}}_{p}-{{\bm{M}}}$ yields:
@@
$\nu_{\mathcal{F}}=\inf_{\mathbf{v}\in\mathbb{E},\norm{v}=1}\mathbf{v}^{\top}\mathcal{F}(\mathbf{v}\mathbf{v}^{\top})v.$ (42)
@@
Therefore, for $\alpha\geq\frac{1}{\nu_{\mathcal{F}}}$ and small enough $\epsilon$, ${{\bm{M}}}^{t}$ expands linearly along each $\mathbf{v}_{i}$.
B.4 Proof of Lemma 4.3
From the definition of $g_{out}$ in Equation 4 and the dominated-convergence theorem, it is straightforward to see that:
$\partial_{{\bm{\omega}}}{\bm{g}}_{\rm out}(Y,{\bm{\omega}}=\mathbf{0},{\bm{I}})=\mathbb{E}[{\bm{Z}}{\bm{Z}}^{\top}-{\bm{I}}|Y].$ (45)
Therefore for any $\mathbf{v}\in\mathbb{R}^{p}$ with $\norm{\mathbf{v}}=1$:
$\mathbf{v}^{\top}\partial_{{\bm{\omega}}}{\bm{g}}_{\rm out}(Y,{\bm{\omega}},{\bm{I}})_{{\bm{\omega}}=\mathbf{0}}\mathbf{v}$ $=\mathbb{E}[(\mathbf{v}^{\top}{\bm{Z}})^{2}-1|Y]$
$=\lim_{d\rightarrow\infty}\mathbb{E}[\rm He_{2}(\langle\mathbf{v}^{\top}{{\bm{{W}}}}^{\star}\mathbf{x}\rangle)|y=Y],$
where the last expectation is w.r.t the joint measure $p(\mathbf{x},y)$ defined in Definition 1.
Now, for any measurable transformation $\mathcal{T}:\mathbb{R}\rightarrow\mathbb{R}$, we have, by the tower law of expectation:
$\mathbb{E}\left[\mathcal{T}(y){\rm He}_{2}(\langle\mathbf{v}^{\top}{{\bm{{W}}}}^{\star},\mathbf{x}\rangle)\right]=\mathbb{E}_{y}\left[\mathcal{T}(y)\mathbb{E}_{\mathbf{x}}\left[\rm He_{2}(\langle\mathbf{v}^{\top}{{\bm{{W}}}}^{\star},\mathbf{x}\rangle)|y\right]\right].$ (46)
For any $\mathbf{v}\in H^{\star}_{E}$, the integrand vanishes asymptotically almost surely over $y$. Therefore, the dominated-convergence theorem implies that for $\mathbf{v}\in H^{\star}_{E}$:
$\lim_{d\rightarrow\infty}\mathbb{E}\left[\mathcal{T}(y){\rm He}_{2}(\langle\mathbf{v}^{\top}{{\bm{{W}}}}^{\star},\mathbf{x}\rangle)\right]=0.$ (47)
B.5 Proof of Theorem 5.1